OpenStudy (anonymous):

find the limit as x approaches negative infinity. -6+(5/x)/6-(1/x^2)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

\[\frac{-6+\frac{5}{x}}{6-\frac{1}{x^2}}\]

OpenStudy (anonymous):

not sure exactly how to read what you wrote, which is why i would be suspicious of any quick answers

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} -6+(5/x)/6-(1/x^2) = 6\]

OpenStudy (anonymous):

I have done this problem before, so I know the answer :D

OpenStudy (anonymous):

then it must be a different problem from the one i wrote.

OpenStudy (anonymous):

Yeah, must be, lol

OpenStudy (blacksteel):

Assuming your function is \[-6 + (5/x)/6 - 1/x^2\], let's quickly simplify it as \[-6 + 5/6x - 1/x^2\] As x approaches negative infinity, the denominators containing x will approach either positive or negative infinity; either way, this will result in the terms approaching 0 (since division by extremely large numbers results in extremely small numbers). Thus, every term but the -6 will be eliminated in the limit, and thus\[\lim_{x \rightarrow -\infty} -6 + 5/6x - 1/x^2 = -6\]

OpenStudy (blacksteel):

(By the way, rgupta was close but left out the negative sign)

OpenStudy (anonymous):

Huh, I'm 98% sure that it is a positive 6..

OpenStudy (anonymous):

well it if is what blacksteel wrote then it is clearly -6 (not 6) with your eyeballs. but i don't know anyone who would write \[(5/x)/6\] which is why i am suspicious

OpenStudy (anonymous):

This is the function: \[ (6+5/(x 6)-1/x^2)\]

OpenStudy (anonymous):

but it is pointless to speculate and whoever posted the question seems not to want to clarify

OpenStudy (anonymous):

It's x6

OpenStudy (blacksteel):

Yeah, that does seem a little bit weird, but it seems like that's what the problem is. If it is the problem you wrote, you can apply the same theory to get \[\lim_{x \rightarrow -\infty} (-6 + 5/x) / (6 - 1/x^2) = -6/6 = 1\]

OpenStudy (blacksteel):

*-1

OpenStudy (anonymous):

Oh, wait. You're right! I accidentally thought that that was a random line on my paper. My bad, sorry. It is -6.

OpenStudy (anonymous):

My teacher put a red spot next to my six, and it covered up my answer >.>

OpenStudy (anonymous):

Assuming your function is −6+(5/x)/6−1/x2

OpenStudy (anonymous):

Jesus, if it the ENTIRE first part over the ENTIRE second part it is -1. I doubt it would be written as -6+(5/x)/6-1/x^2 Do they not think you would be able to handle it if it was written -6+(5/6x)-1/x^2?

OpenStudy (anonymous):

i'd love to know what was actually written on the paper but i guess i never will. @malevolence if it is the second thing you wrote the problem is entirely trivial (which doesn't mean you didn't write it correctly, it just make me wonder why it would be asked)